# Beyond Single-Device Constraints: A System-Level Theoretical Framework for High-Performance Single-Photon Detection at Room Temperature

### v4

@@ 2.2.1 Notations and Basic Assumptions
The parameters governing the system behavior are defined as follows.
As stated, $\eta_{s}$ and $d_{s}$ represent the DE and DCR of $\text{ESPD}_{s}$, respectively. Although $\eta_{s}$ may depend on the photon number distribution, it is approximated as a unified value for the given system for simplicity. The parameters $n_{s+1}$ and $k_{s+1}$ denote the number of effective controlled operations and the decision threshold for $\text{ESPD}_{s+1}$, respectively.
For convenience, all controlled operations $C(F_{i})$ are assumed to be identical and are denoted simply as $C(F)$. They are implemented in a feedforward controlled configuration, such that $n_{s+1}$ represents both the number of effective modules and the total number of physical modules required at level $s+1$. However, the generalized ESPD paradigm remains valid for post-selection controlled gates (see Section 3.1.2 for a brief discussion). Let $p$ be the transmission probability of the signal path through a single module (an (optional) state preparation followed by a $C(F)$). Specifically, $p$ is the probability that the output signal remains non-vacuum given a non-vacuum input of a single non-cascaded module, accounting for loss but irrespective of decoherence (since the state can be re-prepared). Let $P$ and $Q$ denote the probabilities of obtaining a non-vacuum output on the controlled path before detection, given that the input to the corresponding module was non-vacuum or vacuum, respectively. Hence, $Q$ represents the intrinsic error rate of the auxiliary path in the absence of $C(F)$ (i.e., errors from auxiliary state preparation and measurement (SPAM), but not including detection). The parameters $p$, $P$, and $Q$ are approximately assumed to be invariant across all $C(F)$ module operations.
It is crucial to note that $p$, $P$, and $Q$ represent the intrinsic performance metrics of a single, non-cascaded controlled module. These parameters are defined at the component level and remain invariant with respect to the number of cascaded modules. The cumulative effects of loss and error due to the system’s cascaded architecture will be fully incorporated into the derived system equations, such as $\eta_{s+1}$ in Eq.(6).
@@ 2.2.2 Parameter Approximation Assumptions
The assumptions below are employed solely to derive analytical approximations in Section 2.3, as well as the approximation part in Eq. (1) to (3). They are not required for the general theoretical formulation of the ESPD framework nor for the numerical simulations presented in this work. Although these assumptions entail relatively stringent requirements, they are physically motivated by realistic components and will be discussed in detail in Section 3.1.
@@
2.2.3 Formulations
Let $P_{s}$ and $Q_{s}$ denote the probabilities of obtaining a positive report on an auxiliary path given a non-vacuum and vacuum input to a $C(F)$ module, respectively, using $\text{ESPD}_{s}$ as the detector. They can be expressed as:
@@ 2.2.3 Formulations
$P^{\prime}_{s+1}=\eta_{s}+(1-\eta_{s})d_{s}\approx\eta_{s},\ Q^{\prime}_{s+1}=d_{s}$ (3)
The DE and DCR for the $(s+1)$-th level are derived as follows.
For the DE calculation, a non-vacuum input is assumed, and the probability of obtaining at least $k_{s+1}$ positive reports out of $n_{s+1}+1$ total detections is evaluated. If the photon is lost exactly after the $i$-th $C(F)$, the probability of a final positive report is:
$P_{s+1,k_{s+1},i}=$ $(1-Q^{\prime}_{s+1})\sum_{j_{1}+j_{2}=k_{s+1}}^{n_{s+1}}\binom{i}{j_{1}}P_{s+1}^{j_{1}}(1-P_{s+1})^{i-j_{1}}\binom{n-i}{j_{2}}Q_{s+1}^{j_{2}}(1-Q_{s+1})^{n-i-j_{2}}$ (4)
@@
$P_{s+1,k_{s+1}}=$ $P^{\prime}_{s+1}\sum_{j\geq k_{s+1}-1}\binom{n_{s+1}}{j}P_{s+1}^{j}(1-P_{s+1})^{n_{s+1}-j}+(1-P^{\prime}_{s+1})\sum_{j\geq k_{s+1}}\binom{n_{s+1}}{j}P_{s+1}^{j}(1-P_{s+1})^{n_{s+1}-j}$ (5)
$=$ $P^{\prime}_{s+1}\binom{n_{s+1}}{k_{s+1}-1}P_{s+1}^{k-1}(1-P_{s+1})^{n_{s+1}-k_{s+1}+1}+\sum_{j\geq k_{s+1}}\binom{n_{s+1}}{j}P_{s+1}^{j}(1-P_{s+1})^{n_{s+1}-j}$
By combining the above contributions, the DE of $\text{ESPD}_{s+1}$ is given by:
@@
$d_{s+1}=$ $(1-Q^{\prime}_{s+1})\sum_{j\geq k_{s+1}}\binom{n_{s+1}}{j}Q_{s+1}^{j}(1-Q_{s+1})^{n_{s+1}-j}+Q^{\prime}_{s+1}\sum_{j\geq k_{s+1}-1}\binom{n_{s+1}}{j}Q_{s+1}^{j}(1-Q_{s+1})^{n_{s+1}-j}$ (7)
Collectively, the DE and DCR evolution is governed by the non-linear time-variant dynamic system:
$(\eta_{s+1},d_{s+1})=G_{s+1}(\eta_{s},d_{s}),\ (\eta_{0},d_{0})=(\eta,d)$ (8)
where the map $G_{s+1}$ is defined by Eq. (1) to (7).
2.3 Analytical Approximations and Stability Analysis
While Eq. (8) describes the exact dynamics, an analytical solution is generally intractable. However, by employing the approximations in Eq. (1) to (3) together with the assumptions in Section 2.2.2, simplified bounds can be derived to elucidate the system behavior.
The evolution of the DCR is considered first. Eq. (7) can be reformulated as:
$d_{s+1}=$ $Q^{\prime}_{s+1}\binom{n_{s+1}}{k_{s+1}-1}Q_{s+1}^{k-1}(1-Q_{s+1})^{n_{s+1}-k_{s+1}+1}+\sum_{j\geq k_{s+1}}\binom{n_{s+1}}{j}Q_{s+1}^{j}(1-Q_{s+1})^{n_{s+1}-j}$ (9)
@@ 2.3 Analytical Approximations and Stability Analysis
$\frac{d}{dx}f(x)=$ $x^{k-2}[\binom{n}{k-1}a(1-x)^{n-k}[(k-1)(1-x)-x(n-k+1)]$ (11)
$+\sum_{j\geq k}\binom{n}{j}[jx^{j+1-k}(1-x)^{n-j}-x^{j-k+2}(1-x)^{n-j-1}(n-j)]]$
@@
By applying the upper bound from Eq. (2), substituting Eq. (3) into Eq. (9), and assuming $k-1\geq nx$, the following bound is obtained:
$d_{s+1}\leq$ $d_{s}\binom{n_{s+1}}{k_{s+1}-1}(Q+d_{s})^{k_{s+1}-1}(1-Q-d_{s})^{n_{s+1}-k_{s+1}+1}+\sum_{j\geq k_{s+1}}\binom{n_{s+1}}{j}(Q+d_{s})^{j}(1-Q-d_{s})^{n_{s+1}-j}$ (12)
which is notably independent of $\eta_{s}$. Given $Q_{s},d_{s}\ll 1$ and moderate values of $n_{s+1}$ and $k_{s+1}$, it is assumed that $\binom{n_{s+1}}{j}(Q+d_{s})\ll 1$. This yields the approximation:
$d_{s+1}\lesssim d_{s}\binom{n_{s+1}}{k_{s+1}-1}(Q+d_{s})^{k_{s+1}-1}+\binom{n_{s+1}}{k_{s+1}}(Q+d_{s})^{k_{s+1}}=(Q+d_{s})^{k_{s+1}-1}[(d_{s}\binom{n_{s+1}}{k_{s+1}-1}+\binom{n_{s+1}}{k_{s+1}}(Q+d_{s})]$ (13)
This provides an estimation for the DCR at level $s+1$.
@@
$\eta_{s+1}\geq$ $p^{n_{s+1}}[P^{\prime}_{s+1}\sum_{j\geq k_{s+1}-1}\binom{n_{s+1}}{j}P_{s+1}^{j}(1-P_{s+1})^{n_{s+1}-j}+(1-P^{\prime}_{s+1})\sum_{j\geq k_{s+1}}\binom{n_{s+1}}{j}P_{s+1}^{j}(1-P_{s+1})^{n_{s+1}-j}]$ (14)
$=$ $p^{n_{s+1}}[P^{\prime}_{s+1}\binom{n_{s+1}}{k_{s+1}-1}P_{s+1}^{k_{s+1}-1}(1-P_{s+1})^{n_{s+1}-k_{s+1}+1}+\sum_{j\geq k_{s+1}}\binom{n_{s+1}}{j}P_{s+1}^{j}(1-P_{s+1})^{n_{s+1}-j}]$
$\approx$ $p^{n_{s+1}}[\eta_{s}\binom{n_{s+1}}{k_{s+1}-1}(P\eta_{s})^{k_{s+1}-1}(1-P\eta_{s})^{n_{s+1}-k_{s+1}+1}+\sum_{j\geq k_{s+1}}\binom{n_{s+1}}{j}(P\eta_{s})^{j}(1-P\eta_{s})^{n_{s+1}-j}]$
which is independent of $d_{s}$. Therefore, the condition for DE enhancement at the subsequent level is determined by solving the inequality
@@
If $n_{s}$ and $k_{s}$ are constant across levels and the system is stable, the DE of $\text{ESPD}_{s}$ will converge to a root of the left-hand side of Eq. (15).
2.4 Numerical Simulations
While obtaining an exact analytical solution for the non-linear, time-variant dynamic system governing the ESPD in Eq.(8) is mathematically complex, numerical methods provide a robust and sufficient approach for practical design. In this section, numerical simulations of the ESPD paradigm are presented, indicating that substantial performance gains can be achieved even with moderate initial hardware specifications. For example, a three-level ESPD can upgrade a semiconductor SPD to effective performance metrics with DE $93\%$ and DCR $10^{-9}$, comparable to those of state-of-the-art SNSPDs without reliance on superconducting materials.
Simulation Parameters: The performance of the ESPD scheme is evaluated using two distinct baseline SPDs. The first represents an older one, characterized by parameters $(\eta_{0},d_{0})=(59\%,10^{-2})$ [28], while the second represents more modern InGaAs/InP SPD performance, with $(\eta_{0},d_{0})=(27.5\%,10^{-6})$ [29]. Additionally, the controlled-gate parameters are set to $p=0.98$, which can be achieved by commercial devices [38], and $P=0.97$, achievable recently by feedforward C-NOT gate implementations in photonic platforms [39, 40, 41, 42], while, for the auxiliary path, the error rate is set to $Q=0.2\%$, justified by the extremely low SPAM errors (without detections) achievable across specific DOF and the inherent design freedom in selecting the appropriate DOF for the ESPD implementation[43, 38] (see Section 3.1 for parameter discussions). Supplementary simulations with degraded parameter settings are also performed to evaluate robustness in Section 3.1.
Performance Outcomes: Numerical performance metrics across varying configurations and baseline architectures are summarized in Table 2 and 3 with visual representations in Fig. 2. The results indicate that the ESPD scheme can significantly enhance the capabilities of conventional SPDs. Notably, with only three or four levels of enhancement, the detectors achieve performance metrics (DE $93\%$, DCR $10^{-9}$) that are comparable to those of advanced superconducting SPDs.
Table 2: Performance of an ESPD initialized with a SPD from [28], with $\eta_{0}=59\%,\ d_{0}=10^{-2}$.
Para 1 Para 2 Para 3

### v5

@@ 2.2.1 Notations and Basic Assumptions
The evolution of the system is governed by a set of parameters characterizing both detection components and coupling modules.
As defined previously, $\eta_{s}$ and $d_{s}$ represent the effective DE and DCR of the $\text{ESPD}_{s}$ unit, respectively. For analytical tractability, $\eta_{s}$ is treated as a unified value under a given photon-number distribution in the system. The architectural configuration at each level is determined by the number of effective controlled operations $n_{s}$ and the decision threshold $k_{s}$.
We assume that all controlled operations $C(F_{i})$ are identical (denoted as $C(F)$) and implemented in a feed-forward configuration, such that $n_{s}$ represents both the number of effective modules and the total number of physical modules required at level $s$. However, the generalized ESPD paradigm remains valid for post-selection controlled gates (see Section 3.2.2 for a brief discussion). Let $p$ be the transmission probability of the signal path through a single module (an (optional) state preparation followed by a $C(F)$), accounting for incidental losses. Specifically, $p$ is the probability that the output signal remains non-vacuum given a non-vacuum input of a single non-cascaded module, accounting for loss but irrespective of decoherence (since the state can be re-prepared). We further define $P$ and $Q$ as the conditional probabilities of obtaining a non-vacuum output on the auxiliary (controlled) path before detection, given a non-vacuum or vacuum input to the module, respectively. Thus, $Q$ encapsulates the intrinsic error rate of the auxiliary path attributable to state preparation and measurement (SPAM), not including the final detection, in the absence of $C(F)$.
It is crucial to note that $p$, $P$, and $Q$ represent the intrinsic performance metrics of a single, non-cascaded controlled module. These parameters are defined at the component level and remain invariant with respect to the number of cascaded modules. Their cumulative impact on the system’s evolutionary trajectory is fully captured by the recursive relations derived in Section 2.2.3.
@@ 2.2.2 Parameter Approximation Assumptions
While the general formulation remains exact, the following assumptions, which are physically and practically motivated, are introduced to facilitate the analytical approximations in Section 2.3. Such assumptions are not required for the general theoretical formulation of the ESPD framework nor for the numerical simulations and will be discussed in detail in Section 3.2.
@@
2.2.3 Evolutionary Equation Formulations
To describe the state transition from level $s$ to $s+1$, we first evaluate the reporting probabilities on the auxiliary and signal paths. Let $P_{s+1}$ and $Q_{s+1}$ be the probabilities of obtaining a positive report on an auxiliary path using $\text{ESPD}_{s}$ as the detector, given a non-vacuum and vacuum input to the $C(F)$ module, respectively:
@@ 2.2.3 Evolutionary Equation Formulations
$P^{\prime}_{s+1}=\eta_{s}+(1-\eta_{s})d_{s}\approx\eta_{s},\quad Q^{\prime}_{s+1}=d_{s}$ (3)
The performance of $\text{ESPD}_{s+1}$ is determined by the collective response of $n_{s+1}+1$ detectors. For the DE calculation, a non-vacuum input is assumed, and the probability of obtaining at least $k_{s+1}$ positive reports out of $n_{s+1}+1$ total detections is evaluated. If the photon is lost exactly after the $i$-th $C(F)$, the probability of a final positive report is:
$P_{s+1,k_{s+1},i}$ (4)
$=$ $(1-Q^{\prime}_{s+1})\sum_{j_{1}+j_{2}=k_{s+1}}^{n_{s+1}}\binom{i}{j_{1}}P_{s+1}^{j_{1}}(1-P_{s+1})^{i-j_{1}}\binom{n-i}{j_{2}}Q_{s+1}^{j_{2}}(1-Q_{s+1})^{n-i-j_{2}}$
@@
$P_{s+1,k_{s+1}}=$ $P^{\prime}_{s+1}\sum_{j\geq k_{s+1}-1}\binom{n_{s+1}}{j}P_{s+1}^{j}(1-P_{s+1})^{n_{s+1}-j}$ (5)
$\qquad\qquad\qquad\ +(1-P^{\prime}_{s+1})\sum_{j\geq k_{s+1}}\binom{n_{s+1}}{j}P_{s+1}^{j}(1-P_{s+1})^{n_{s+1}-j}$
$=$ $P^{\prime}_{s+1}\binom{n_{s+1}}{k_{s+1}-1}P_{s+1}^{k-1}(1-P_{s+1})^{n_{s+1}-k_{s+1}+1}$
$\qquad\qquad\qquad\qquad\qquad\quad+\sum_{j\geq k_{s+1}}\binom{n_{s+1}}{j}P_{s+1}^{j}(1-P_{s+1})^{n_{s+1}-j}$
The total effective DE at level $s+1$ is then the weighted sum:
@@
$d_{s+1}=(1-Q^{\prime}_{s+1})\sum_{j\geq k_{s+1}}$ $\binom{n_{s+1}}{j}Q_{s+1}^{j}(1-Q_{s+1})^{n_{s+1}-j}$ (7)
$+Q^{\prime}_{s+1}\sum_{j\geq k_{s+1}-1}\binom{n_{s+1}}{j}Q_{s+1}^{j}(1-Q_{s+1})^{n_{s+1}-j}$
In summary, the transition of performance metrics follows a non-linear dynamical system:
$(\eta_{s+1},d_{s+1})=G_{s+1}(\eta_{s},d_{s}),\quad(\eta_{0},d_{0})=(\eta,d)$ (8)
where the map $G_{s+1}$ is defined by the recursive operators in Eqs. (1) to (7).
The ESPD architecture thus induces a nonlinear dynamics whose fixed points, stability, and basins of attraction determine the asymptotic detection performance. Particularly, the fixed point $(\eta_{\infty},d_{\infty})$ represents the fundamental physical limit of the ESPD framework under given architectural parameters, while the basins of attraction define the range of hardware-level imperfections that can be effectively mitigated to achieve such enhancement.
2.3 Analytical Approximations Analysis
While Eq. (8) describes the exact dynamics of the ESPD framework, an explicit analytical solution for the iterated map is generally intractable. However, by employing the approximations in Eqs. (1) to (3) alongside the physically-motivated assumptions in Section 2.2.2, we can derive simplified bounds to elucidate the system’s evolutionary behavior.
Consider the evolution of DCR first. Eq. (7) can be reformulated as:
$d_{s+1}=Q^{\prime}_{s+1}\binom{n_{s+1}}{k_{s+1}-1}Q_{s+1}^{k-1}(1$- $Q_{s+1})^{n_{s+1}-k_{s+1}+1}$ (9)
$+\sum_{j\geq k_{s+1}}\binom{n_{s+1}}{j}Q_{s+1}^{j}(1-Q_{s+1})^{n_{s+1}-j}$
To analyze the monotonicity of this transition, we introduce the following proposition.
@@ 2.3 Analytical Approximations Analysis
$\frac{d}{dx}f(x)$ (11)
$=$ $x^{k-2}[\binom{n}{k-1}a(1-x)^{n-k}[(k-1)(1-x)-x(n-k+1)]$
$\qquad\qquad\qquad\quad+\sum_{j\geq k}\binom{n}{j}[jx^{j+1-k}(1-x)^{n-j}-x^{j-k+2}(1-x)^{n-j-1}(n-j)]]$
@@
By applying the upper bound from Eq. (2), substituting Eq. (3) into Eq. (9), and assuming $k-1\geq nx$, we obtain the following bound for the DCR evolution:
$d_{s+1}\leq d_{s}\binom{n_{s+1}}{k_{s+1}-1}(Q+d_{s})$ ${}^{k_{s+1}-1}(1-Q-d_{s})^{n_{s+1}-k_{s+1}+1}$ (12)
$+\sum_{j\geq k_{s+1}}\binom{n_{s+1}}{j}(Q+d_{s})^{j}(1-Q-d_{s})^{n_{s+1}-j}$
Notably, this bound is independent of the DE $\eta_{s}$. Given $Q,d_{s}\ll 1$ and moderate values for the architectural parameters $(n,k)$, we can assume $\binom{n_{s+1}}{j}(Q+d_{s})\ll 1$. This yields the approximation:
$d_{s+1}$ $\lesssim d_{s}\binom{n_{s+1}}{k_{s+1}-1}(Q+d_{s})^{k_{s+1}-1}+\binom{n_{s+1}}{k_{s+1}}(Q+d_{s})^{k_{s+1}}$ (13)
$=(Q+d_{s})^{k_{s+1}-1}[(d_{s}\binom{n_{s+1}}{k_{s+1}-1}+\binom{n_{s+1}}{k_{s+1}}(Q+d_{s})]$
This expression provides a practical estimation for the DCR suppression at the iterative level $s+1$.
@@
$\eta_{s+1}\geq$ $p^{n_{s+1}}[P^{\prime}_{s+1}\sum_{j\geq k_{s+1}-1}\binom{n_{s+1}}{j}P_{s+1}^{j}(1-P_{s+1})^{n_{s+1}-j}$ (14)
$\qquad\qquad\qquad\qquad+(1-P^{\prime}_{s+1})\sum_{j\geq k_{s+1}}\binom{n_{s+1}}{j}P_{s+1}^{j}(1-P_{s+1})^{n_{s+1}-j}]$
$=$ $p^{n_{s+1}}[P^{\prime}_{s+1}\binom{n_{s+1}}{k_{s+1}-1}P_{s+1}^{k_{s+1}-1}(1-P_{s+1})^{n_{s+1}-k_{s+1}+1}$
$\qquad\qquad\qquad\qquad\qquad\qquad\ +\sum_{j\geq k_{s+1}}\binom{n_{s+1}}{j}P_{s+1}^{j}(1-P_{s+1})^{n_{s+1}-j}]$
$\approx$ $p^{n_{s+1}}[\eta_{s}\binom{n_{s+1}}{k_{s+1}-1}(P\eta_{s})^{k_{s+1}-1}(1-P\eta_{s})^{n_{s+1}-k_{s+1}+1}$
$\qquad\qquad\qquad\qquad\qquad\qquad\ +\sum_{j\geq k_{s+1}}\binom{n_{s+1}}{j}(P\eta_{s})^{j}(1-P\eta_{s})^{n_{s+1}-j}]$
This bound is independent of $d_{s}$. Hence, the condition for DE enhancement at the subsequent level (namely $\eta_{s+1}>\eta_{s}$) is determined by the inequality:
@@
If $n_{s}$ and $k_{s}$ are constant across levels and the system is stable, the DE of the $\text{ESPD}_{s}$ will (approximately) converge to a stable fixed point corresponding to a root of the left-hand side of Eq. (15).
2.4 Monte Carlo Steady-State Analysis and Numerical Simulations
While obtaining an exact analytical solution for the non-linear dynamic system governing the ESPD in Eq. (8) is mathematically complex, numerical methods provide a robust approach for stability analysis and practical design. In this section, we present the Monte Carlo analysis to characterize the system’s evolutionary behavior, including the location of stable fixed points and their basins of attraction, and provide numerical simulations under practically motivated device parameters. The results indicate that substantial performance gains can be achieved even with moderate initial hardware specifications. For instance, a three-level ESPD can elevate a conventional semiconductor SPD to effective performance metrics (DE $93\%$ and DCR $10^{-9}$), matching state-of-the-art SNSPDs, but without requiring cryogenic infrastructure.
Simulation Parameters: The parameter settings are physically motivated by commercial specifications or experimental reports. The controlled-gate parameters are set to $p=0.98$ and $P=0.97$, values recently demonstrated in optical platforms and feedforward C-NOT gate implementations on photonic platforms [38, 39, 40, 41, 42]. For the auxiliary path, the error rate is set to $Q=0.2\%$, justified by the extremely low SPAM errors achievable across specific DOF and the inherent design freedom in selecting the appropriate DOF for the ESPD implementation [43, 42]. Detailed discussions on parameters and robustness under degraded settings are provided in Section 3.2.
[…]
Performance Evolution on Legacy SPD: We specifically examine the evolutionary trajectory of the ESPD initialized with a legacy SPD from [28] (DE $\approx 59\%$, DCR $\approx 10^{-2}$). Table 2 details the step-by-step evolutions, with visual representations in Fig. 5. These results illustrate that the ESPD scheme significantly transcends the intrinsic limitations of conventional SPDs. Notably, within only three to four levels of iterative enhancement, the effective metrics (DE $93\%$, DCR $10^{-9}$) become comparable to advanced superconducting SPDs, consistent with the stability predicted by our steady-state analysis.
(a) Steady-state DE $\eta_{\infty}$ vs. initial DE and DCR $(\eta_{0},d_{0})$.
(b) Steady-state DCR $d_{\infty}$ vs. initial DE and DCR $(\eta_{0},d_{0})$.
Figure 2: Monte Carlo steady-state analysis of DE and DCR of the ESPD framework for $(n_{s},k_{s})=(8,4)$ at each level $s$.
(a) Steady-state DE $\eta_{\infty}$ vs. initial DE and DCR $(\eta_{0},d_{0})$.
(b) Steady-state DCR $d_{\infty}$ vs. initial DE and DCR $(\eta_{0},d_{0})$.
Figure 3: Monte Carlo steady-state analysis of DE and DCR of the ESPD framework for $(n_{s},k_{s})=(6,3)$ at each level $s$.
(a) Steady-state DE $\eta_{\infty}$ vs. initial DE and DCR $(\eta_{0},d_{0})$.
(b) Steady-state DCR $d_{\infty}$ vs. initial DE and DCR $(\eta_{0},d_{0})$.
Figure 4: Monte Carlo steady-state analysis of DE and DCR of the ESPD framework for $(n_{s},k_{s})=(4,2)$ at each level $s$.
Table 2: Performance of an ESPD initialized with an SPD from [28], with $\eta_{0}=59\%,\ d_{0}=10^{-2}$.
Para 1: $(n,k)=(4,2)$ Para 2: $(n,k)=(6,3)$ Para 3: $(n,k)=(8,4)$